OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 2 7 12 17 y = __x + __

OpenStudy (anonymous):

@fallenangelorchid @quin100 @quesia @Write0The0Right0For0Wrong0Doing @whydoihavetosignup1 @emo__gamer101 @Goku-Kai @just_one_last_goodbye @hugsnotughs @jordanloveangel @CausticSyndicalist @Michele_Laino @gelow @gwenny_78 @Warrior678 @kyley_bear @oscar55773 @undeadknight26

OpenStudy (anonymous):

WHAT?

OpenStudy (anonymous):

sorry for caps

OpenStudy (anonymous):

I had to type the graph

OpenStudy (anonymous):

For every change of x by 1, y increases by 5. So the gradient of the line is 5. To find the y-intercept: y = mx + c 2 = 5(1) + c 2 = 5 + c c = -3

OpenStudy (michele_laino):

since you are searching for this equation,: \[y=ax+b\] if I set x=1,and y=2 into my formula, and I get:\[2=a+b\]

OpenStudy (michele_laino):

similarly if I use the second point, namely if I set x=2 and y=7 into my formula, I get: \[7=2a+b\]

OpenStudy (anonymous):

Hey when you finish helping @LDShadowLady , could you guys help me out???

OpenStudy (michele_laino):

Now we have got a system of two equations, namely: \[2=a+b,7=2a+b\] please solve that equations for a and b

OpenStudy (michele_laino):

ok! @kyley_bear

OpenStudy (michele_laino):

please, do you know how to solve that system? @LDShadowLady

OpenStudy (michele_laino):

@LDShadowLady

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! Then try, please!

OpenStudy (anonymous):

is it y = 5x + -3

OpenStudy (anonymous):

its alright

OpenStudy (michele_laino):

that's right! @LDShadowLady y=5x-3

OpenStudy (anonymous):

ok thanks